# remarkRemark

### v1

@@ 3 Preliminaries and Assumptions
${\mathbb{X}}^{B}_{l}:=\bigotimes_{s\in B}{\mathbb{X}}^{s}_{l},\qquad{\mathbb{X}}^{s}_{l}:=\operatorname{conv}\left(\{y\in{\mathbb{R}}^{n_{l}}\;|\;\exists\;v\in U_{l-1}\;s.t.\;y=\lambda_{l}(x,v),\;{\mathbf{x}}\in{\mathbb{X}}^{s}_{l-1}\}\right),$
where ${\mathbb{X}}^{s}_{0}=\{x^{s}\}$ and ’$\operatorname*{conv}$’ denotes the convex hull. Correspondingly, to emphasize the dependency of a feature map upon a fixed set of trainable parameters, we set ${\mathbf{x}}_{l}({\mathbf{u}})\in{\mathbb{X}}_{l}^{B}$ to represent the entirety of feature maps at a given layer, generated using the trainable parameters ${\mathbf{u}}\in U_{ad}$. In line with this notation, the following compact notation is introduced for each layer function corresponding to a given data batch
@@
    The function $\Phi_{B}\in C^{1}({\mathbb{X}}_{L}^{B},{\mathbb{R}})$ is bounded from below by zero and there exists a constant $C>0$ such that for all ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{X}}_{L}^{B}$ it holds
    $\|\Phi_{B}({\mathbf{x}})-\Phi_{B}({\mathbf{y}})\|+\|\nabla\Phi_{B}({\mathbf{x}})-\nabla\Phi_{B}({\mathbf{y}})\|\leq C\|{\mathbf{x}}-{\mathbf{y}}\|.$
@@
Assumption A1 on the terminal loss function is satisfied by state-of-the-art mean squared loss used for regression tasks and the cross-entropy loss commonly applied in classification with CNNs [19]. To consider the applicability of the Assumptions A2 to the standard machine learning framework, we give the following Proposition for non-residual structures.
@@
For ease of notation, we consider the single sample case $|B|=1$ and include the bias function into the definition of ${\cal W}_{l}$ by adjusting the network’s states correspondingly. Thus, without loss of generality, we can suppose ${\cal B}\equiv 0$. Using elementary Lipschitz estimates and the boundedness of the Jacobian $\partial f_{l}$ by assumption, we have that for all $({\mathbf{x}},u),({\mathbf{y}},v)\in{\mathbb{X}}^{B}_{l}\times U_{l}$ there exists a constant $C>0$ such that
(7) $\|\partial_{x}\Lambda_{l}({\mathbf{x}},u)-\partial_{x}\Lambda_{l}({\mathbf{y}},v)\|_{M}\leq\|{\cal W}_{l}(u_{l})\|^{2}_{M}\|{\mathbf{x}}-{\mathbf{y}}\|+C\|{\cal W}_{l}(u_{l})\|^{2}_{M}\|{\mathbf{y}}\|\|u-v\|.$
A similar reasoning applies to obtain a Lipschitz estimate for the layer functions $\Lambda_{l}$. While the compactness of $U_{l}$ together with the Lipchitz property of the mapping ${\cal W}_{l}$ guarantee the boundedness of $\|{\cal W}_{l}(u_{l})\|_{M}$, it remains to prove the boundedness of ${\mathbb{X}}_{l}^{B}$. By the composite structure of (1), we have ${\mathbf{x}}_{l+1}({\mathbf{u}})=f_{l}({\cal W}_{l}(u_{l}){\mathbf{x}}_{l}({\mathbf{u}}))$. Then by the Lipschitz property of $f_{l}$ for all $l=0,\ldots,L-1$, there exists a constant $C>0$ such that the following inequality holds:
@@
$\|{\mathbf{x}}_{l+1}({\mathbf{u}})\|\leq C\|x^{s}\|+\sum_{l=0}^{L-1}\|f_{l}(0)\|,$
@@
By Proposition 3.1, we can ensure that Assumption A2 holds in our machine learning framework if $\rho>0$ and a lower semi-continuous and coercive regularizer ${\cal R}$ is used. This combination, coupled with the boundedness from below of $\Phi_{B}$, guarantees that the corresponding level set $\{{\mathbf{u}}\;|\;J_{B}({\mathbf{u}})\leq C\}\subseteq U_{ad}$ is compact, for every $C>0$ chosen sufficiently large. Consequently, the existence of a minimizer for (4) is assured, and we can accommodate the Lipschitz property required in A2 for the subsequent results by selecting $U_{ad}$ corresponding to an appropriate level set.
The case where $\rho=0$ is also well-posed. In fact, the existence of a minimizer for (4) can be guaranteed through a so-called max-norm regularization technique, which was initially introduced for CNNs in [31]. In our context, this technique can be modeled by defining ${\cal W}_{l}$ as the composition of a projection operation onto a given compact and convex set followed by a linear transformation. This adjustment not only preserves the Lipschitz property of the weight function but also ensures the boundedness of ${\mathbb{X}}^{B}$, and thus the validity of A2. Additionally, due to the structure of our neural network, the regularity assumption in A2 regarding the states is satisfied for classical sigmoidal activation functions and average pooling operations. Although functions like ReLU and max-pooling operations are not included, we remark that these functions can be smoothly approximated up to any desired degree of accuracy by using functions such as Softplus for ReLU and a weighted Softmax function instead of max-pooling. For a comprehensive overview of commonly used activation functions, we refer to [32]. Consequently, we expect that our results will hold at least asymptotically for ReLU-based networks, ensuring the practical application of our method to a variety of cases. This claim is supported by numerical evidence presented in Section 8.
@@ 4 The maximum principle for CNNs
The state-of-the-art approach in machine learning for solving the problem (4) is to apply a so-called backpropagation technique, introduced by Rumelhart et al. [33]. This scheme relies on gradient-based optimization methods and is in its essence a combination of the chain rule and classical gradient descent [19, 20]. Commonly for this approach, it is assumed that the empirical risk function $J_{B}$ is differentiable. Various derivations of the backpropagation methodology were introduced, leading to multiple variants of the original algorithm. For our framework, the approach by LeCun [2] is especially outstanding, as it motivates backpropagation through optimal control-based techniques by invoking the variational principles and the Lagrange formalism to formulate first-order necessary conditions for supervised learning problems of the form (4). Under suitable differentiability assumptions, these first-order necessary conditions are formulated as follows:
@@
    (10) $\nabla_{{\mathbf{u}}_{l}}J_{B}({\mathbf{u}}^{*}):=\nabla_{{\mathbf{u}}_{l}}\left(R_{l}({\mathbf{u}}^{*})-{\mathbf{p}}_{l+1}({\mathbf{u}}^{*})\cdot\Lambda_{l}({\mathbf{x}}({\mathbf{u}}^{*}),u^{*}_{l})\right)=0,$
@@
(11) $\mathcal{H}^{B}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),u_{l}):={\mathbf{p}}_{l+1}({\mathbf{u}})\cdot\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})-R_{l}(u_{l}).$
With this auxiliary function at hand the system (8)-(10) can equivalently be defined in terms of the HP functions gradients, leading to the observation that for each layer, (11) is maximized along the optimal trajectory generated by ${\mathbf{u}}^{*}$. The original statement of the PMP is due to Pontryagin et al. [17, 18]. One decade after this first formulation of the maximum principle, Halkin [16] derived a version of the PMP for discrete-time optimal control problems, which was shortly after further generalised by Holtzmann [34]. An additional generalization of the PMP was presented by Zhan et al. [35] who formulated a unified theory for the maximum principle and weakened the regularity assumptions given by Halkin for the discrete-time case. Based upon [35], in our framework, the statement of the discrete PMP is as follows:
@@
In its discrete-time version and opposed to the original statement of the PMP in continuous-time, one assumption is the convexity of the sets $R_{l}(U_{l})$ and $\Lambda_{l}({\mathbf{x}},U_{l})$ for any fixed ${\mathbf{x}}\in{\mathbb{X}}_{l}^{B}$. We remark that due to the affine structure of the pre-activation mechanism, we have that in our framework, $\Lambda_{l}({\mathbf{x}},U_{l})$ is convex for any fixed ${\mathbf{x}}\in{\mathbb{X}}_{l}^{B}$ if the set $U_{l}$ is convex. This condition also implies the convexity of $R_{l}(U_{l})$ in the standard CNN framework with a continuous regularizer; see e.g. [14]. It is important to remark that these convexity assumptions are not required for the construction of our PMP-based algorithm. The maximality condition (12) allows the formulation of backpropagation algorithms in cases where no gradient or subgradient is available. That is, for example, where the set of admissible parameters is discrete [14, 26, 27] or the regularizer is only lower semi-continuous; for the latter, see Section 6.
@@ 5 Generalized training based on the PMP
Having a characterization of the optimal feature maps and backpropagation variables in the form of (8) - (9) together with the maximality condition stated in the discrete PMP, the backpropagation algorithm can be motivated as a forward-backwards sweep method. This class of indirect methods aims to derive an optimal solution ${\mathbf{u}}^{*}$ by first processing forward through the structure (8) utilizing some previously chosen ${\mathbf{u}}^{(k)}$, followed by determining the terminal cost functions gradients to then propagate them backwards using (9). After all the variables ${\mathbf{x}}_{l}({\mathbf{u}}^{(k)})$ and ${\mathbf{p}}_{l}({\mathbf{u}}^{(k)})$ have been determined, the network’s parameters are updated based upon the condition (10). Corresponding to the classical backpropagation approach, this update is commonly determined with one step of a gradient descent-based strategy. After this update the same process is then repeated iteratively, using the new parameters ${\mathbf{u}}^{(k+1)}$. While this classical backpropagation approach is widely applied in machine learning, it is generally not applicable in cases where no gradient information with respect to the trainable parameters is available. That is, if the set of admissible parameters is discrete or a discontinuous regularizer is considered. However, considering instead a maximization of the underlying HP function gives rise to a backpropagation approach which includes gradient-based updating steps as a special case and accommodates training even without available gradient information. The basis for this generalization is the so-called method of successive approximations (MSA) which was introduced by Krylov and Chernous’ko in the early 1960s [21], see also [22]. In our discrete-time setting, and in the spirit of the discrete PMP, the MSA is updating the parameters of a CNN by a maximization step on the discrete Hamiltonian as follows
@@
While fast convergence with these MSA schemes was achieved in special cases, the method appeared less robust with respect to parameter changes and initialization for general non-linear control systems. As a result, in its basic form, the MSA method is inappropriate for a direct application in the framework of machine learning due to its lack of stability. However, in the last decades, several, robust variants of the original MSA for continuous-time optimal control problems have been proposed [12, 23, 24, 25] including modifications specifically developed for application to the framework of supervised learning with neural networks [14, 13]. The sequential quadratic Hamiltonian (SQH) approach for machine learning [13], which is inspired by its continuous-time equivalent [25, 36, 37], enforces the convergence of the basic MSA-scheme by adjusting the Hamiltonian maximization step (13) using an augmented Hamiltonian approach [23, 38]. That is, for updating the trainable parameters at each iteration, the classical HP function in (13) is replaced by the following augmented HP function
@@
where the augmentation strength $\epsilon>0$ is adjusted suitably at each iteration, to guarantee a reduction of the loss functional value. The use of an augmented HP function is motivated by the observation that a decrease in the full batch loss is achieved by increasing the value of the HP function by simultaneously keeping the parameter changes at each iteration moderate; see Theorem 7.3. While the application of the SQH scheme to the framework of fully-connected ResNet-like architectures is comparably straightforward [13], much work is required for its extension to the framework of CNNs. In fact, extending the SQH method to CNNs presents significant challenges. One major difficulty involves modeling complexities introduced by convolution operations with multiple channels and pooling layers. This together with the consideration of high-dimensional inputs (e.g. high-resolution RGB images), which leads to an increased computational complexity of the training procedure, requires the formulation of special techniques. A particular challenge in this context is that the resulting increased inference times and memory usage make the exact maximization of the augmented Hamiltonian practically unfeasible in most cases. To face this challenge, we can replace the exact maximization of the augmented Hamiltonian by the less strong requirement that each update $u_{l}^{(k+1)}\in U_{l}$ for all $l=1,\ldots,L-1$, is such that the following inequality holds:
@@
While this modification makes an updating step more feasible numerically, the increased inference times and possible hardware restrictions make processing of the complete available data non-economical in practice. Motivated by this challenge, we propose a further relaxation of the updating step by invoking a stochastic approach. The basis for this approach is splitting the available data set $B$ into so-called mini-batches of size $M<|B|$ and training at each iteration on one of these randomly sampled subsets. That is, at each iteration, prior to executing (8) - (9) a batch of data is randomly sampled from
(16) $B(M):=\{B_{k}\subset 2^{B}\;|\;|B_{k}|=M\leq|B|\},$
where in the following, we denote the corresponding set of indices for a mini-batch $B_{k}$ by $I_{B_{k}}$. The forward-backward procedure then is executed on this mini-batch of data and the trainable parameters are adjusted based a mini-batch specific augmented HP function, where the Hamiltonian is replaced by its mini-batch variant as follows:
(17) ${\cal H}^{B_{k}}_{l}({\mathbf{x}},{\mathbf{p}},u):=\sum_{s\in I_{B}}\left[p^{s}\cdot\lambda_{l}(x^{s},u)\right]\chi_{I_{B_{k}}}(s)-R_{l}(u),$
where $\chi_{I_{B_{k}}}:I_{B}\rightarrow\{0,1\}$ defines the indicator function. As a basis for this modification we establish (17) as an unbiased estimator for the full-batch HP function (11).
@@
Let $B_{k}\in B(M)$ be drawn uniformly at random. Then for each fixed ${\mathbf{u}},{\mathbf{w}}\in U_{ad}$ and all $l=0,\ldots,L-1$ it holds
$\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l})\right]={\cal H}^{B}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l}).$
@@
Due to the fixed parameter choice, we can without loss of generality for this proof, assume that $\rho=0$. Since $B_{k}\in B(M)$ is drawn uniformly at random, due to the fundamental bridge between expectation and probability, we have
@@
We define for each $s\in I_{B}$, $q^{s}_{l}({\mathbf{u}})=Mp^{s}_{l}({\mathbf{u}})$ and achieve for the HP functions realizations the following form
(19) ${\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l})=\frac{1}{|B|}\sum_{s\in I_{B}}\left[q^{s}_{l+1}({\mathbf{u}})\cdot\lambda_{l}(x^{s}_{l}({\mathbf{u}}),w_{l})\right]\frac{|B|}{M}\chi_{I_{B_{k}}}(s).$
Using the result (18), together with the linearity of the expected value yields
$\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l})\right]$ $=\frac{1}{|B|}\sum_{s\in I_{B}}\left[q^{s}_{l+1}({\mathbf{u}})\cdot\lambda_{l}(x^{s}_{l}({\mathbf{u}}),u_{l})\right]\frac{|B|}{M}\mathbb{E}\left[\chi_{I_{B_{k}}}\right]$
@@
$\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}^{*}),{\mathbf{p}}_{l+1}({\mathbf{u}}^{*}),u^{*}_{l})\right]\geq\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}^{*}),{\mathbf{p}}_{l+1}({\mathbf{u}}^{*}),u)\right],\qquad u\in U_{l}.$
This result, together with the law of large numbers motivates replacing the full-batch Hamiltonian in the augmented HP function (15) by its mini-batch counterpart (17) and thus gives rise to the following batch SQH (bSQH) scheme:
@@
       while $J_{B_{k}}\left({\mathbf{w}}\right)-J_{B_{k}}\left({\mathbf{u}}^{(k)}\right)>-\eta{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{w}}-{\mathbf{u}}^{(k)}|\kern-1.07639pt|\kern-1.07639pt|}$ do
@@
by choosing the minimal $j\in{\mathbb{N}}$ such that a decrease in the mini-batch loss is achieved. It is apparent that the choice of $\hat{\epsilon}_{k}$ has a decisive impact on the number of required maximization steps and thus on the efficiency of the scheme, due to the increased inference complexity. While the basic SQH scheme uses a backtracking approach as follows

### v2

@@ 3 Preliminaries and Assumptions
${\mathbb{X}}^{B}_{l}:=\bigotimes_{s\in I_{B}}{\mathbb{X}}^{s}_{l},\qquad{\mathbb{X}}^{s}_{l}:=\operatorname{conv}\left(\{y\in{\mathbb{R}}^{n_{l}}\;|\;\exists\;v\in U_{l-1}\;s.t.\;y=\lambda_{l-1}(x,v),\;x\in{\mathbb{X}}^{s}_{l-1}\}\right),$
where ${\mathbb{X}}^{s}_{0}=\{x^{s}\}$ and $\operatorname*{conv}$’ denotes the convex hull operator. Correspondingly, to emphasize the dependency of a feature map upon a fixed set of trainable parameters, we denote with ${\mathbf{x}}_{l}({\mathbf{u}})\in{\mathbb{X}}_{l}^{B}$ the entirety of feature maps at a given layer, generated using the trainable parameters ${\mathbf{u}}\in U_{ad}$. Based thereon, the following compact notation is introduced for each layer function corresponding to a given data batch
@@
    The function $\Phi_{B}\in C^{1}({\mathbb{X}}_{L}^{B},{\mathbb{R}})$ is bounded from below and there exists a constant $C>0$ such that for all ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{X}}_{L}^{B}$ it holds
    $\|\nabla\Phi_{B}({\mathbf{x}})-\nabla\Phi_{B}({\mathbf{y}})\|\leq C\|{\mathbf{x}}-{\mathbf{y}}\|.$
@@
The regularity assumption in A1 on the terminal loss function is satisfied by the state-of-the-art mean squared loss used for regression tasks and the cross-entropy loss commonly applied in classification with CNNs [1]. To consider the applicability of the Lipschitz assumptions in A1-A2 to the standard neural network setting, we prove the following Proposition for non-residual structures.
@@
For ease of notation, we consider the single sample case $|B|=1$ and merge the bias function into the definition of ${\cal W}_{l}$ by adjusting the network’s states correspondingly. Thus, without loss of generality, we can suppose ${\cal B}\equiv 0$. Using elementary Lipschitz estimates and the boundedness of the Jacobian $\partial f_{l}$ by assumption, we have that for all $({\mathbf{x}},u),({\mathbf{y}},v)\in{\mathbb{X}}^{B}_{l}\times U_{l}$ there exists a constant $C>0$ such that
(7) $\|\partial_{x}\Lambda_{l}({\mathbf{x}},u)-\partial_{x}\Lambda_{l}({\mathbf{y}},v)\|_{M}\leq C\left(\|{\cal W}_{l}(u)\|^{2}_{M}\|{\mathbf{x}}-{\mathbf{y}}\|+\|{\cal W}_{l}(u_{l})\|_{M}\|{\mathbf{y}}\|\|u-v\|\right).$
A similar reasoning applies to obtain a Lipschitz estimate for the layer functions $\Lambda_{l}$. The compactness of $U_{l}$ together with the Lipschitz property of the mapping ${\cal W}_{l}$ guarantees the boundedness of $\|{\cal W}_{l}(u)\|_{M}$ and it remains to prove the boundedness of each ${\mathbb{X}}_{l}^{B}$. By definition, we have ${\mathbf{x}}_{l+1}({\mathbf{u}})=f_{l}({\cal W}_{l}(u_{l}){\mathbf{x}}_{l}({\mathbf{u}}))$ and by the Lipschitz property of $f_{l}$ for all $l=0,\ldots,L-1$, there exists a constant $C>0$ such that the following inequality holds:
@@
$\|{\mathbf{x}}_{l+1}({\mathbf{u}})\|\leq C\left(\|x^{s}\|+\sum_{l=0}^{L-1}\|f_{l}(0)\|\right),$
@@
[…]
The case where $\rho=0$ is also well-posed. In fact, the existence of a minimizer for (4) can be guaranteed through a so-called max-norm regularization technique, which was initially introduced for CNNs in [39]. In our context, this technique can be modelled by defining ${\cal W}_{l}$ as the composition of a projection operation onto a given compact and convex set, followed by a linear transformation. This adjustment not only preserves the Lipschitz property of the weight function but also ensures the boundedness of ${\mathbb{X}}^{B}$, and thus the validity of A1-A2. Additionally, due to the structure of our neural network, the regularity assumption in A2 regarding the states is satisfied for classical sigmoidal activation functions and average pooling operations. We note that, even though functions like ReLU and max-pooling operations are not included, these functions can be smoothly approximated up to any desired degree of accuracy by, e.g. Softplus for ReLU and a weighted Softmax function for max-pooling. For a comprehensive overview of commonly used activation functions, we refer to [14]. Consequently, we expect that our results will hold at least asymptotically for ReLU-based networks, ensuring the practical application of our method to a variety of cases. This is supported by numerical evidence presented in Section 8.
@@ 4 The maximum principle for CNNs
The state-of-the-art approach in machine learning for solving the problem (4) is to apply a so-called backpropagation technique, introduced by Rumelhart et al. [35]. This scheme relies on gradient-based optimization methods and is, in its essence, a combination of the chain rule and classical gradient descent [1, 33]. Commonly, for this approach, it is assumed that the empirical risk function $J_{B}$ is differentiable. Various derivations of the backpropagation methodology have been introduced, leading to multiple variants of the original algorithm. For our framework, the approach by LeCun [27] is especially outstanding as it motivates backpropagation through optimal control-based techniques by invoking variational principles and the Lagrange formalism to formulate first-order necessary conditions for supervised learning problems of the form (4). Under suitable differentiability assumptions, these first-order necessary conditions are formulated as follows:
@@
    (10) $\nabla_{u_{l}}J_{B}({\mathbf{u}}^{*}):=\nabla_{u_{l}}\left(\rho R_{l}({\mathbf{u}}^{*})-{\mathbf{p}}_{l+1}({\mathbf{u}}^{*})\cdot\Lambda_{l}({\mathbf{x}}({\mathbf{u}}^{*}),u^{*}_{l})\right)=0,$
@@
(11) $\mathcal{H}^{B}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),u_{l}):={\mathbf{p}}_{l+1}({\mathbf{u}})\cdot\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})-\rho R_{l}(u_{l}).$
With this auxiliary function at hand the system (8)-(10) can equivalently be defined in terms of the HP functions gradients, leading to the observation that for each layer, (11) is maximized along the optimal trajectory generated by $\mathbf{u}$$^{*}$. The original statement of the PMP is due to Pontryagin et al. [34, 4]. One decade after this first formulation of the maximum principle, Halkin [20] derived a version of the PMP for discrete-time optimal control problems, which was shortly after further generalized by Holtzmann [23]. An additional generalization of the PMP was presented by Zhan et al. [44] who formulated a unified theory for the maximum principle and weakened the regularity assumptions given by Halkin for the discrete-time case. Based upon [44], in our framework, the statement of the discrete PMP is as follows:
@@
In its discrete-time version and opposed to the original statement of the PMP in continuous-time, a main assumption is the convexity of the sets $R_{l}(U_{l})$ and $\Lambda_{l}({\mathbf{x}},U_{l})$ for any fixed ${\mathbf{x}}\in{\mathbb{X}}_{l}^{B}$. We remark that due to the affine structure of the pre-activation mechanism, we have that in our framework, $\Lambda_{l}({\mathbf{x}},U_{l})$ is convex for any fixed ${\mathbf{x}}\in{\mathbb{X}}_{l}^{B}$ if the set $U_{l}$ is convex. This condition also implies the convexity of $R_{l}(U_{l})$ in the standard CNN framework with a continuous regularizer; see e.g. [29]. It is important to remark that these convexity assumptions are not required for the construction of our PMP-based algorithm. The maximality condition (12) allows the formulation of backpropagation algorithms in cases where no gradient or subgradient is available. That is, for example, where the set of admissible parameters is discrete [29, 2, 3] or the regularizer is only lower semi-continuous; for the latter, see Section 6.
@@ 5 Generalized training based on the PMP
Having a characterization of the optimal feature maps and backpropagation variables in the form of (8) - (9) together with the maximality condition stated in the discrete PMP, the backpropagation algorithm can be motivated as a forward-backwards sweep method. This class of indirect methods aims to derive an optimal solution $\mathbf{u}$$^{*}$ by first processing forward through the structure (8) utilizing some previously chosen ${\mathbf{u}}^{(k)}$, followed by determining the terminal cost functions gradients, to then propagate them backwards using (9). After all the variables ${\mathbf{x}}_{l}({\mathbf{u}}^{(k)})$ and ${\mathbf{p}}_{l}({\mathbf{u}}^{(k)})$ have been determined, the network’s parameters are updated based upon the condition (10). Corresponding to the classical backpropagation approach, this update is commonly determined with one step of a gradient descent-based strategy. The process is then repeated iteratively, using the new parameters ${\mathbf{u}}^{(k+1)}$. While this classical backpropagation approach is widely applied in machine learning, it is generally not applicable in cases where no gradient or subgradient information with respect to the trainable parameters is available. That is, if the set of admissible parameters is discrete or a discontinuous regularizer is used. However, considering a condition similar to (12) for updating the network’s parameters gives rise to a backpropagation approach which includes gradient-based updating steps as a special case and accommodates training even without available gradient information. The basis for this generalization is the so-called method of successive approximations (MSA), which was introduced by Krylov and Chernous’ko in the early 1960s [26]; see also [24]. In our discrete-time setting, and in the spirit of the discrete PMP, the MSA is updating the parameters of a CNN by a maximization step on the discrete Hamiltonian as follows
@@
While fast convergence with these MSA schemes was achieved in special cases, the method appeared less robust with respect to parameter changes and initialization for general nonlinear control systems. As a result, in its basic form, the MSA method is inappropriate for direct application in the framework of machine learning due to its lack of stability. However, in the last decades, several robust variants of the original MSA for continuous-time optimal control problems have been proposed [28, 37, 10, 6], including modifications specifically developed for application to the framework of supervised learning with neural networks [29, 22]. The sequential quadratic Hamiltonian (SQH) approach for machine learning [22], which is inspired by its continuous-time equivalent [6, 5, 7], enforces the convergence of the basic MSA-scheme by adjusting the Hamiltonian maximization step (13) using an augmented Hamiltonian approach [37, 38]. That is, for updating the trainable parameters at each iteration, the classical HP function in (13) is replaced with the following augmented HP function
@@
where the augmentation strength $\epsilon>0$ is adjusted suitably at each iteration, to guarantee a reduction of the loss functional value. The use of an augmented HP function is motivated by the observation that a decrease in the full batch loss is achieved by increasing the value of the HP function while simultaneously keeping the parameter changes at each iteration moderate; see Theorem 7.3. While the application of the SQH scheme to the framework of fully-connected ResNet-like architectures is comparably straightforward [22], much work is required for its extension to the framework of CNNs. In fact, extending the SQH method to CNNs presents significant challenges. One major difficulty is the exact maximization of the augmented HP function, which may be computationally infeasible due to large-scale input data (e.g., high-resolution RGB images) and the complexity of modern CNN architectures with millions of trainable parameters and deep hierarchical structures. To face this challenge, we replace the exact maximization of the augmented Hamiltonian with the relaxed requirement that for all $l=1,\ldots,L-1$, each update $u_{l}^{(k+1)}\in U_{l}$ must only satisfy the following:
@@
While this makes the updating step more feasible numerically, processing the complete dataset at once can still lead to longer inference times, increased memory requirements and a costly, potentially prohibitive training process. This motivates a further relaxation of the updating step by invoking a stochastic approach. The basis for this approach is splitting the available dataset $B$ into so-called mini-batches of size $M\leq|B|$ and training at each iteration on one of these randomly sampled subsets. That is, at each iteration, before executing (8) - (9), a batch of data is randomly sampled from
(16) $B(M):=\{B_{k}\subset 2^{B}\;|\;|B_{k}|=M\},$
where in the following, we denote the corresponding set of indices for a mini-batch $B_{k}$ by $I_{B_{k}}$. The forward-backwards procedure then is executed on this mini-batch of data and the trainable parameters are adjusted based on a mini-batch-specific augmented HP function, where the Hamiltonian is replaced with its mini-batch variant as follows:
(17) ${\cal H}^{B_{k}}_{l}({\mathbf{x}},{\mathbf{p}},u):=\sum_{s\in I_{B}}\left[p^{s}\cdot\lambda_{l}(x^{s},u)\right]\chi_{I_{B_{k}}}(s)-\rho R_{l}(u),$
where $\chi_{I_{B_{k}}}:I_{B}\rightarrow\{0,1\}$ defines the indicator function. As a basis for this modification, we establish (17) as an unbiased estimator for the full-batch HP function (11).
@@
Let $B_{k}\in B(M)$ be drawn uniformly at random and define ${\mathbf{q}}_{l}({\mathbf{u}}):=\frac{|B|}{M}{\mathbf{p}}_{l}({\mathbf{u}})$ for all $l\in\{0,\ldots,L\}$. Then for each fixed ${\mathbf{u}},{\mathbf{w}}\in U_{ad}$ and all $l=0,\ldots,L-1$ it holds
$\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{q}}_{l+1}({\mathbf{u}}),w_{l})\right]={\cal H}^{B}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l}).$
@@
We assume, without loss of generality for this proof, that $\rho=0$, as the regularizer is independent of each mini-batch. Since $B_{k}\in B(M)$ is drawn uniformly at random and due to the fundamental bridge between expectation and probability, we have
@@
We reformulate the realizations of the mini-batch HP function as follows
${\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{q}}_{l+1}({\mathbf{u}}),w_{l})=\frac{M}{|B|}\sum_{s\in I_{B}}\left[q^{s}_{l+1}({\mathbf{u}})\cdot\lambda_{l}(x^{s}_{l}({\mathbf{u}}),w_{l})\right]\frac{|B|}{M}\chi_{I_{B_{k}}}(s).$
Thus, using the result (18), together with the linearity of the expected value, yields
$\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}),{\mathbf{q}}_{l+1}({\mathbf{u}}),w_{l})\right]$ $=\sum_{s\in I_{B}}\left[p^{s}_{l+1}({\mathbf{u}})\cdot\lambda_{l}(x^{s}_{l}({\mathbf{u}}),w_{l})\right]\frac{|B|}{M}\mathbb{E}\left[\chi_{I_{B_{k}}}\right]$
@@
$\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}^{*}),{\mathbf{q}}_{l+1}({\mathbf{u}}^{*}),u^{*}_{l})\right]\geq\mathbb{E}\left[{\cal H}^{B_{k}}_{l}({\mathbf{x}}_{l}({\mathbf{u}}^{*}),{\mathbf{q}}_{l+1}({\mathbf{u}}^{*}),u)\right],\qquad u\in U_{l}.$
This result, together with the law of large numbers, motivates replacing the full-batch Hamiltonian in the augmented HP function (15) by its mini-batch counterpart (17) and thus gives rise to the following batch SQH (bSQH) scheme:
@@
[…]
@@
by choosing the minimal $j\in{\mathbb{N}}$ such that a decrease in the mini-batch loss is achieved. It is apparent that the choice of $\hat{\epsilon}_{k}$ has a decisive impact on the number of required inference steps and thus on the efficiency of the scheme. While the basic SQH scheme uses a backtracking approach as follows